OpenStudy (anonymous):

lim as x tends to infinity of (1+3/x)^x

OpenStudy (tkhunny):

You may wish to introduce a logarithm.

OpenStudy (kirbykirby):

Recall the definition of the important limit: \[ \lim_{x\rightarrow \infty} \left(1+\frac{a}{x} \right)^x=e^a\]

OpenStudy (kirbykirby):

Oh yes maybe it is too easy to use that.. lol back to first principles I guess

OpenStudy (kirbykirby):

Yes the logarithm trick is probably one of the easier ways to prove that limit. What you can do is let \[y=\lim_{x\rightarrow\infty}\left( 1+\frac{3}{x}\right)^x\] then take the log of both sides : \[\begin{aligned} \log y &= \log \left[ \lim_{x\rightarrow\infty}\left( 1+\frac{3}{x}\right)^x\right]\\ &=\lim_{x\rightarrow\infty}\log\left[\left( 1+\frac{3}{x}\right)^x\right]\\& = \lim_{x\rightarrow\infty}x\log\left( 1+\frac{3}{x}\right)\\ & =\lim_{x\rightarrow\infty}\frac{\log\left( 1+\frac{3}{x}\right)}{\frac{1}{x}}\end{aligned}\] This should give you a good start